Find the value of m so that


lim (sin 2x + m sin 3x)/ x^3


x→0


is finite.

1
Expert's answer
2021-11-22T20:20:55-0500

"\begin{aligned}
&\operatorname{lim}_{x \rightarrow 0} \frac{\sin 2 x+m \sin 3x}{x^{3}} \\
&lim_{x \rightarrow 0} \frac{[2 \sin x\cos x+m(3\sin x-4\sin^3x)]}{x^{3}} \\
&lim_{x \rightarrow 0} \frac{\sin x[2 \cos x+m(3-4\sin^2x)]}{x^{3}} \\
&lim_{x \rightarrow 0} \frac{[2 \cos x+m(3-4\sin^2x)]}{x^{2}} \\
&\{\because lim_{x \rightarrow 0} \frac{\sin x}{x}=1 \}\\

\end{aligned}"

As "x \rightarrow 0" , denominator tends to , so the numerator also tends to 0.

Thus, "lim_{x \rightarrow 0} \ [2 \cos x+m(3-4\sin^2x)]=0\\
2+3m=0\\
m=\frac{-2}{3}"